### v1

# Fairness And Performance In Harmony: Data Debiasing Is All You Need

Abstract
Fairness in both machine learning (ML) predictions and human decisions is critical, with ML models prone to algorithmic and data bias, and human decisions affected by subjectivity and cognitive bias. This study investigates fairness using a real-world university admission dataset with 870 profiles, leveraging three ML models: XGB, Bi-LSTM, and KNN. Textual features are encoded with BERT embeddings. For individual fairness, we assess decision consistency among experts with varied backgrounds and ML models, using a consistency score. Results show ML models outperform humans in fairness by 14.08% to 18.79%. For group fairness, we propose a gender-debiasing pipeline and demonstrate its efficacy in removing gender-specific language without compromising prediction performance. Post-debiasing, all models maintain or improve their classification accuracy, validating the hypothesis that fairness and performance can coexist. Our findings highlight ML’s potential to enhance fairness in admissions while maintaining high accuracy, advocating a hybrid approach combining human judgement and ML models.
Index Terms:
Fairness, Responsible AI, Data Debiasing
**footnotetext: Corresponding author: j@forth.ai
I Introduction
Fairness has emerged as a critical concern in terms of both the predictive outputs of Machine Learning (ML) models and the decisions made by human experts [1, 2, 3, 4]. In the context of ML models, challenges in fairness often stem from data biases and algorithmic limitations, which can propagate or exacerbate existing inequities. Conversely, fairness in human decision-making is inherently influenced by subjective judgment and cognitive biases, making it vulnerable to inconsistency and error. These two sources of unfairness necessitate rigorous investigation into both computational and human-centric frameworks to mitigate bias and ensure equitable outcomes.
To investigate fairness, this work utilizes a real-world university admission dataset with 870 unique profiles and three effective ML models, namely, Extreme Gradient Boosting (XGB), Bi-directional Long Short-Term Memory (Bi-LSTM) and K-Nearest Neighbours (KNN). The textual features of the profiles are encoded into high-dimensional embeddings using the Bidirectional Encoder Representations from Transformers (BERT), enabling the extraction of rich contextual representations for downstream predictive tasks. In our case, we focus on the important but challenging task of admission offer decision making.
We begin by examining individual fairness in offer decisions for a heterogeneous group of applicants, characterized by diverse backgrounds, leadership qualities, and academic achievements. Using a standardized set of features for comparison, we assess individual fairness by measuring decision consistency through a consistency score [5], which quantifies how similarly applicants with comparable profiles are treated by human experts or ML models. Our results reveal that both XGBoost and Bi-LSTM outperform human experts significantly, achieving higher consistency scores by a margin of 14.08% to 18.79%. These findings suggest that ML models demonstrate superior reliability in ensuring fair treatment of similar applications.
Furthermore, we assess group fairness by analyzing the effects of gender debiasing on classification performance and consistency across different models and decision-making stages. To address potential biases, we propose a debiasing pipeline designed to eliminate gender-specific language from the input features. The effectiveness of this approach is demonstrated through a statistical analysis, confirming its capability to mitigate gender bias while preserving the integrity of classification outcomes.
Subsequently, we examine the effects of data debiasing on prediction performance by employing three empirically effective machine learning models to perform classification tasks on the dataset, both before and after debiasing. Experimental results across all three models validate our hypothesis, demonstrating that predictions either maintain or improve in accuracy following the debiasing process.
TABLE I: Data Format and Sample Entry
Field Format Sample Entry
GCEA [GCEA] School:{school}, UAS:{score}; Grades:{subject grade pairs} [GCEA]: School:HCI, UAS:90.0; Grades:H1 PROJECT WORK A, H1 GENERAL PAPER A, H2 ECONOMICS A, H2 CHEMISTRY A, …
GCEO [GCEO] {subject-grade pairs} [GCEO] HIGHER CHINESE B3, ENGLISH A1, ELEMENTARY MATHEMATICS A2, …
Leadership Multiple Entries: [[Leadership] {activity}, Level:{role}, Year:{year}, Category:{type}, Participation:{level}]. [Leadership] Mind Sports Club, Level:Captain, Year:2023, Category:Sports, Participation:Executive Committee
PIQ [PIQ{n}]{essay text}, where n = 1,2,3,4,5 [PIQ1] My first visit to the university in Primary 5 for a Learning Journey left me with a deep impression…
OfferType Offered / Not Offered Offered
Contributions
In summary, the main contributions of this work are as follows:
  1. 1.
    We propose robust pipelines for identifying and mitigating biases in machine learning models and statistically validate their effectiveness.
  2. 2.
    We experimentally demonstrate that debiasing data not only leads to comparable results but often enhances model performance, challenging the assumption that debiasing sacrifices accuracy.
  3. 3.
    We conducted a rigorous empirical study on individual and group fairness in a real-world university admission setting, revealing that ML models can outperform human experts in consistency while maintaining or surpassing them in making fair and accurate decisions.
II Related Work
II-A Algorithmic Fairness
The notion of individual fairness, introduced  [6], emphasizes treating similar individuals similarly. This definition of individual fairness is closely related to differential privacy [7]. Specifically, for every pair of individuals, the Lipschitz condition holds. Individual fairness, however, does not imply group fairness, which focuses on ensuring that different groups (often defined by protected attributes such as race, gender, or age) are treated fairly. What is fair or not depends on the worldview held [8]. Fairness measures consistent with the worldview can often be derived from confusion matrices.
An interesting development in the field of computer science is the increasing use of large language models (LLMs) to automate decision tasks. As discussed in [9], human-like biases may be picked up by LLMs. It is important to investigate whether these biases exist in real life applications and whether or not these biases may be mitigated.
Prior works also investigated methods to mitigate bias to enhance fairness in various aspects. For instance, [10] proposed strategies through robust data preprocessing to achieve fair and accurate predictive modelling in the education domain. [5] proposed debiasing algorithms for fair representation learning. [11] provided a comprehensive study on auditing techniques for human expertise.
II-B Fairness in University Admissions
There has been numerous works studying various aspects of fairness in admissions but many of these are based on smaller-scale qualitative studies [12], general admissions policies [13] or specific aspects like standardized tests [14]. While these studies provide interesting insights, their methodologies often require significant effort to replicate. Furthermore, the small-scale user groups or the limited geographic focus of these works constrain the generalizability of their findings to broader contexts. Other studies, such as [15], focus on comparing human perceptions of fairness in AI-driven decision-making with those of human decision-making.
Puranik et al. [16] proposed the Fair-Greedy selection policy where participants selection from under-represented groups reinforces their participation in future rounds. Fair-Greedy balances between two objectives of greedy (maximize scores of selected applicants) and fairness (minimize deviations of the proportion of under-represented group).
Bhattacharya et al. [17] conducted an empirical investigation into biases in university admissions arising from applicants’ socioeconomic factors. Their study utilized a threshold-crossing model, integrated with admission data, to analyze disparities in admission thresholds across different demographic groups and identify the directionality of these differences.
II-C Differences with Earlier Work
While existing literature offers valuable insights into algorithmic fairness and fairness in education, limited attention has been given to addressing biases present in text-derived data, which are inherently prone to subjectivity. To the best of our knowledge, this study is the first to demonstrate an effective approach for mitigating such biases to enhance individual fairness. Furthermore, we experimentally prove that debiasing, counterintuitively, not only preserves but often improves model prediction performance.
Figure 1: Proposed pipelines to assess (a)Individual Fairness and (b)Group Fairness
III Data Processing
III-A Dataset
We use the university admission dataset introduced in [18], which consists of applicant profiles from the 2024 admission cycle. Table I contains a summary of the fields, formats and sample entries of the original data. The dataset includes a total of 870 unique profiles that come with four features: General Certificate of Education Advanced Level (GCEA), General Certificate of Education Ordinary Level (GCEO), Personal Insight Questions (PIQ) and Leadership Experience. We concatenate all content and form a Combined document.
III-B Gender Debiasing
The debiasing pipeline focuses on removing gender-specific language from the PIQ and Leadership features while preserving other meaningful content. We maintained a comprehensive list of gender-specific terms, including pronouns (e.g., ”he”, ”she”), titles (e.g., ”Mr.”, ”Mrs.”), and gender-specific roles (e.g., ”son”, ”daughter”). These terms were systematically removed using a combination of spaCy-based tokenization and regular expression pattern matching.
To assess the effectiveness of debiasing, we perform a binary classification using a pretrained BERT model (bert-base-uncased) and compare the distributions before and after debiasing. The profiles are classified as male or female with a threshold of 0.8, otherwise as unknown.
Table II summarize the gender inference results before and after debiasing. The debiasing intervention demonstrated remarkable efficacy, with the system’s gender predictions showing substantial changes among 870 profiles.
Specifically, the proportion of definitive gender assignments decreased dramatically, with female classifications falling by 83.6% (from 46.7% to 7.7%) and male classifications declining by 75.1% (from 43.6% to 10.9%). As a result, the ”Unknown” category takes up 81% the profiles, rising from 9.6% before debiasing. These results strongly suggest that the debiasing procedure successfully reduces the system’s reliance on gender-based features while maintaining balanced treatment across gender categories.
III-C Feature Embedding
To prepare the data for fairness assessment, we generate contextual embeddings using BERT (bert-base-uncased). For each feature, we generated BERT embeddings using the following process: the text was tokenized and padded to a maximum length of 512 tokens, then processed through BERT to obtain contextual embeddings. We extracted the [CLS] token embedding (768-dimensional vector) as the representation for each feature. The individual embeddings were then concatenated to create a combined representation (3840-dimensional vector) for each profile, preserving the contextual information. For both original and debiased data, the final output included individual embeddings for each feature and embeddings for the concatenated feature.
III-D Human Decisions
We use three human decision points, such as Shortlisting (SL), Admission Recommendation (AR) and Offer (OF), as targets to assess prediction accuracy and consistency. Specifically, we mapped the categorical labels to binary values, where ”Shortlisted”, ”Recommended” and ”Offered” corresponds to 1, and 0 otherwise, respectively.
III-E Data Splitting
To ensure consistent evaluation, we implemented a stratified data splitting strategy using sklearn’s $train_{t}est_{s}plit$ function with a fixed random seed of 42. The data was split into 80% training, 10% validation, and 10% test data. The same split indices were applied to both original and debiased data to maintain consistency in our comparative analysis.
TABLE II: Gender Inference Results Before and After Debiasing
Category Before After Change (%)
Female 524 (46.7%) 86 (7.7%) -83.6%
Male 489 (43.6%) 122 (10.9%) -75.1%
Unknown 108 (9.6%) 913 (81.4%) +745.4%
Total Samples 870
TABLE III: Hyperparameter Search Spaces and Optimal Values
Model Hyperparameter Search Space Optimal
XGB n_estimators randint(100, 1000) 415
max_depth randint(3, 15) 11
learning_rate uniform(0.001, 0.3) 0.014
subsample uniform(0.7, 0.3) 0.61
colsample_bytree uniform(0.7, 0.3) 0.78
reg_alpha uniform(0, 2) 1.13
reg_lambda uniform(0, 4) 1.54
Bi-LSTM hidden_dim [128, 256, 512, 1024, 2048] 1024
dropout [0.1, 0.2, 0.3, 0.4, 0.5, 0.6] 0.6
learning_rate [1e-4, 1e-3, 3e-4, 3e-3, 5e-4, 5e-3] 1e-3
batch_size [1, 4, 8, 16, 32, 64] 1
KNN k [3, 4, 5, 6, 7] 5
similarity [”cosine”, ”L2”] L2
IV Individual Fairness
IV-A Formulation
Individual fairness is grounded in the principle that similar individuals should receive similar treatment in decision-making processes. Following [5], we operationalize this concept through a consistency metric that evaluates how similarly a classifier treats applicants with comparable profiles. Figure 1(a) demonstrates the proposed pipeline of assessing individual fairness.
Formally, given a set of N applicants, we define the consistency score $\mathcal{C}$ of a classifier as:
$\mathcal{C}=1-\frac{1}{N}\sum_{i=1}^{N}|\hat{y}_{i}-\frac{1}{k}\sum_{j\in knn(i)}\hat{y}_{j}|$ (1)
where $\hat{y}_{i}$ represents the classifier’s prediction for applicant $i$, and $\frac{1}{k}\sum_{j\in knn(i)}\hat{y}_{j}$ computes the average of predictions among the k-nearest neighbors (i.e. most similar profiles), of applicant $i$. This formulation yields a score between 0 and 1, where higher values indicate greater consistency in treating similar cases similarly. The scores effectively reflect the fairness consistency by measuring mean absolute differences between a classifier’s predictions to the target profiles and the average predictions made for most similar profiles.
IV-B Research Questions
The rest of the section discusses our investigation on the following Research Questions:
  • RQ1: Can we detect cognitive bias in human decisions that reduce individual fairness?
  • RQ2: Are AI models more consistent in making decisions than human?
  • RQ3: Can AI models outperform human in making admission decisions?
TABLE IV: Individual Fairness: classification performance and fairness consistency. Best results are bolted.
Model Precision Recall F1-Score Accuracy Consistency (AR) Consistency (OF)
Human Decisions
Shortlist(SL) 0.8464 0.8418 0.8156 0.8155 - -
Adm. Rec. (AR) 0.8011 0.8193 0.8321 0.8087 0.5632 -
Offer (OF) - - - - - 0.6023
ML Models
KNN (k=5) 0.6886 0.6707 0.6897 0.6897 - -
XGB 0.7902 0.7859 0.7878 0.7931 0.7521 0.7477
Bi-LSTM 0.8291 0.8178 0.8176 0.8276 0.8073 0.7797
IV-C Methodology
Our evaluation framework examines individual fairness across the entire admission pipeline by analyzing three human decision points, such as shortlisting (SL), admission recommendations (AR), and final offers (OF), alongside two model-based classifiers (XGB and BiLSTM) and a retrieval-based k Nearest Neighbor (KNN) classifier. The foundation of our methodology lies in the construction of a feature-reranked similarity matrix, which provides a robust measure of application similarity based on embedded features. This similarity matrix enables the identification of nearest neighbors for consistency evaluation.
IV-C1 Implementation
For model training and evaluation, we choose three different approaches, namely, traditional classifiers, neural networks and retrieval-based classifiers. For each category, we choose an empirically effective model from each category, such as XGB, BiLSTM, and KNN.
The XGB classifier is implemented using the xgboost library. The BiLSTM model was implemented using PyTorch, featuring a bidirectional LSTM layer followed by linear layers with ReLU activation and LogSoftmax output. For the KNN approach, we utilized the FAISS library for similarity search with in-built memory-efficient processing predictions in batches.
IV-C2 Hyperparameters
We use the validation set to perform hyperparameter search with randomised scheme. Each run takes 20 epochs with a patience of 5 (i.e. 5 epochs without improvement). Best sets are determined using average accuracy score. Table III summarises the search space and optimal set for each classifiers.
IV-C3 Metrics
For each decision point and model, we compute both traditional classification metrics (precision, recall, F1-score, and accuracy) and consistency score. The consistency evaluation uses k=5 nearest neighbors, chosen to balance between local similarity and statistical stability. We maintain consistent feature representations across all models by using the same embedded feature space, ensuring fair comparisons between human decisions and machine learning predictions. This approach allows us to systematically assess how different decision-making approaches align with the principle of treating similar applications similarly, while also measuring their overall classification performance.
IV-D Experimental Results
Table IV summarizes the experimental results for classification performance and individual fairness consistency.
IV-D1 Human Decision vs Model Performance
Human Decisions: The shortlisting decisions (SL) made by human evaluators achieve the highest precision among all decision points and models, with a precision of 84.64%. This high precision indicates strong reliability in correctly identifying qualified candidates at the initial stage of the admission process. The admission recommendation (AR) stage exhibits slightly lower but still robust performance, with an accuracy of 80.87% and an F1-score of 83.21%. These metrics suggest consistent and effective decision-making during the early phases of candidate evaluation process.
Machine Learning Models: Among the machine learning models, the Bi-LSTM model demonstrates superior performance, achieving an F1-score of 81.76% and an accuracy of 82.76%. These results closely match those of human decisions, with the Bi-LSTM model attaining a comparable F1-score and even surpassing human decisions in accuracy. This indicates that the Bi-LSTM model is effective in replicating human-like decision-making patterns while offering enhanced accuracy.
In contrast, the KNN model with $k=5$ shows significantly lower performance, with both F1-score and accuracy at 68.97%. This suggests that a non-parametric retrieval model like KNN is less capable of capturing the complex patterns inherent in human decision-making, highlighting the importance of model selection in predictive tasks.
IV-D2 Fairness Consistency
In terms of fairness consistency, we observe notable variations between human decisions and machine learning model predictions. The consistency score for the admission recommendations (AR) stage is 56.32%, while the final offer decisions (OF) exhibit a slightly higher consistency at 60.23%. These figures suggest that human decision-making maintains moderate levels of consistency, but there remains considerable room for improvement in ensuring that similar candidates receive similar outcomes.
Conversely, both machine learning models demonstrate substantially higher consistency scores compared to human decisions. The XGB model achieves consistency scores of 75.21% for AR and 74.77% for OF stages, indicating a significant improvement over human decision-making in terms of fairness consistency. The Bi-LSTM model performs even better, attaining the highest consistency scores of 80.73% for AR and 77.97% for OF. This highlights the potential of machine learning models to provide more consistent and fair evaluations, reducing variability in decision outcomes for similar applicants.
IV-E Analysis and Interpretations
IV-E1 RQ1: Identifying Cognitive bias
While early stages of human decisions, such as Shortlisting (SL) and Admission Recommendations (AR) demonstrate strong classification performance i.e. over 8% in F1-Score and accuracy, their consistency scores are notably lower, with AR’s being 56.32% and OF’s being 60.23%. This disparity between performance metrics and consistency scores suggests potential cognitive biases in the evaluation process. Furthermore, the gradual decrease in consistency from OF to AR stages, i.e. from 60.23% to 56.32%, indicates that biases might become more pronounced in later stages of the admission process, possibly due to increased complexity in decision criteria while maintaining high consistency among different review panels.
IV-E2 RQ2: Consistency comparison between AI and Human
The consistency analysis reveals a substantial gap between human and machine learning approaches. Both ML models significantly outperform human decisions in terms of consistency, with Bi-LSTM achieving scores of 80.73% and 77.97% for AR and OF respectively, compared to human scores of 56.32% (AR) and 60.23% (OF). This represents an improvement of approximately 24% in consistency for AR decisions and 18% for OF decisions. The XGB model also demonstrates superior consistency with 75.21% for AR and 74.77% for OF, suggesting that machine learning approaches inherently provide more standardized evaluation patterns. This substantial difference in consistency scores highlights the potential role of ML models in reducing decision variability.
IV-E3 RQ3: Accuracy comparison between AI and Human
The performance comparison between human decisions and ML models yields nuanced insights. While human shortlisting achieves the highest precision of 84.64%, the Bi-LSTM model demonstrates comparable overall performance with the highest accuracy of 82.76% and F1-score of 81.76%. Furthermore, the Bi-LSTM’s performance metrics closely match or exceed human decision points while maintaining significantly higher consistency scores. The XGB model, while showing lower performance metrics with an F1-score of 78.78%, still maintains higher consistency than human decisions, suggesting a potential trade-off between performance and consistency that varies across different ML architectures.
Discussion
These findings have several important implications for admission processes:
  • The high consistency of ML models, particularly Bi-LSTM (AR: 80.73%, OF: 77.97%), coupled with their strong performance metrics, suggests they could serve as effective standardization tools in the admission process.
  • The superior precision of human shortlisting (84.64%) combined with the high consistency of ML models indicates that a hybrid approach might be optimal, leveraging the strengths of both human and machine evaluation.
  • The progressive decline in human consistency scores across decision stages suggests a need for standardized decision support tools, particularly in later stages of the admission process.
  • The performance gap between simpler (KNN) and more sophisticated (Bi-LSTM) ML models indicates that architectural choices significantly impact both accuracy and consistency metrics.
In summary, the experimental results demonstrate that while human decisions maintain high classification performance, they show lower consistency compared to ML approaches. These findings suggest that ML models could serve as valuable decision support tools, particularly in maintaining consistency across the admission process while preserving high accuracy standards. The results advocate for a carefully designed hybrid approach that combines human expertise with ML-driven consistency checks to optimize both performance and fairness in admission decisions.
V Group Fairness
V-A Formulation
Group fairness in machine learning systems addresses systematic biases that can disadvantage certain demographic groups in model predictions. We formulate group fairness using accuracy parity, which ensures equal accuracy among groups. In other words, assuming a binary protected attribute $a$, accuracy parity requires:
$\Delta_{Acc}=|P(\hat{Y}=Y|a=0)-P(\hat{Y}=Y|a=1)|=0,$
where $\Delta_{Acc}$ is the difference in accuracy between groups and $\hat{Y}$ represents a ML classifier’s predictions. In the literature, the opposite of accuracy parity is known as disparate mistreatment [19].
Let $D_{original}=\{(X_{original},Y)\}$ be our original dataset, where $X_{original}$ are the original input features and $Y$ are binary labels (offered / not offered); Let $D_{debiased}=\{(X_{debiased},Y)\}$ be the debiased dataset, where $X_{debiased}$ are the features after a debiasing features.
Formally, we define the discrimination score $\Delta_{P}$ as:
$\Delta_{P}=|\Delta_{Acc,D_{original}}-\Delta_{Acc,D_{debiased}}|$
V-B Research Questions
The rest of the section discusses our investigation on the following Research Questions:
  • RQ4: How to identify gender bias in human decisions?
  • RQ5: Does debiasing features sacrifice classification performance?
  • RQ6: Is there a correlation between model complexity and ability to handle biased features?
V-C Methodology
Figure 1(b) illustrates the proposed pipeline for assessing group fairness. The original data first undergoes debiasing steps to removes gender-related clues from the features (details in section III-B. Both original and debiased datasets are then transformed into contextual embeddings using a pretrained BERT language model (bert-base-uncased), with each feature encoded as a 768-d vector.
Subsequently, the embeddings are used as input features for classification. Separate classifiers such as XGB and Bi-LSTM are trained to predict the offer decisions, while KNN being non-parametric directly performs prediction without training.
To evaluate the impact of debiasing, we compute the classification accuracy, consistency and discrimination scores between the original and debiased datasets, using the same hyperparameters as described in section IV to ensure consistency in the analysis.
TABLE V: Group Fairness: accuracy and consistency comparisons among ML classifiers using original and debiased data.
Model Data Accuracy Consistency $\Delta_{P}$
XGB Original 0.8046 0.7303 0.0
Debiased 0.8046 0.7266
BiLSTM Original 0.8276 0.7567 0.0195
Debiased 0.8471 0.7673
KNN Original 0.6322 - 0.0345
Debiased 0.6667
V-D Experimental Results
Table V presents the performance metrics for the XGB, KNN, and BiLSTM models on both the original and debiased datasets. Accuracy represents the proportion of correct predictions, while Consistency measures the stability of model predictions across different data samples. The discrimination score $\Delta_{P}$ quantifies the demographic parity by measuring the absolute difference in model accuracies between original and debiased data.
The BiLSTM model demonstrates superior performance, with its accuracy improving from 82.76% to 84.71% after applying data debiasing techniques. Similarly, its consistency score increases from 75.67% to 76.73%, indicating enhanced robustness. The model exhibits a discrimination score $\Delta_{P}$ of 0.0195, suggesting moderate sensitivity to gender-based patterns in the data. The XGB model maintains a stable accuracy of 80.46% across both datasets, with a minimal decrease in consistency from 73.03% to 72.66%. Notably, its discrimination score $\Delta_{P}$ is 0, indicating complete invariance to the debiasing process. The KNN model, serving as a baseline, shows a notable improvement in accuracy from 63.22% to 66.67% when trained on debiased data, accompanied by the highest discrimination score $\Delta_{P}$ of 0.0345, demonstrating significant sensitivity to bias in the original data.
V-E Analysis and Interpretations
V-E1 RQ4: Assessing Gender Bias in Decision-Making Patterns
The high consistency scores observed in the BiLSTM and XGB models suggest the presence of systematic patterns in the decision-making data. The significant performance gap between the KNN model and the more complex models indicates that these patterns are intricate and require advanced algorithms to model effectively. This complexity may be attributed to underlying biases, such as gender bias, influencing human decisions. The varying discrimination scores ($\Delta_{P}$) across models provide quantitative evidence of these systematic biases, with simpler models showing greater sensitivity to their presence.
V-E2 RQ5: Impact of Debiasing on Model Performance
The results challenge conventional assumptions about debiasing potentially degrading model performance. Instead, the BiLSTM model’s improved accuracy (from 82.76% to 84.71%) and consistency (from 75.67% to 76.73%), coupled with a moderate $\Delta_{P}$ of 0.0195, imply that debiasing can enhance the model’s ability to generalize by removing noise associated with biased features. The KNN model’s substantial improvement in accuracy and high $\Delta_{P}$ of 0.0345 further supports this notion, highlighting that simpler models benefit significantly from data that more accurately reflects unbiased patterns.
V-E3 RQ6: Relationship Between Model Complexity and Bias Handling
The varying responses to debiased data among different models underscore the relationship between model complexity and bias handling. This relationship is quantitatively demonstrated through the progressive increase in discrimination scores: XGB ($\Delta_{P}=0$), BiLSTM ($\Delta_{P}=0.0195$), and KNN ($\Delta_{P}=0.0345$). The KNN model’s high $\Delta_{P}$ and sensitivity to biased data results in greater performance gains when such biases are mitigated. In contrast, the XGB model’s $\Delta_{P}$ of 0 suggests that its ensemble approach inherently manages bias more effectively. The BiLSTM model, despite its complexity, still shows a moderate $\Delta_{P}$ and benefits from debiasing, indicating that even sophisticated models are not immune to the effects of biased data.
Discussion
These findings contribute to a deeper understanding of how model architectures interact with biased data and the effectiveness of debiasing techniques. The discrimination scores $\Delta_{P}$ provide a quantitative measure of each model’s sensitivity to bias, revealing a clear inverse relationship between model complexity and bias sensitivity. The consistent improvement or stability in model performance upon debiasing suggests that addressing biases aligns with performance optimization goals. The varying $\Delta_{P}$ values across model architectures emphasize the importance of considering both model complexity and potential data biases when developing predictive systems. Consequently, practitioners should carefully evaluate their choice of model architecture and implement appropriate debiasing techniques to ensure both ethical integrity and high performance.
VI Limitations and Future Works
While this study provides valuable insights into fairness in decision-making processes, one notable limitation is that we only assessed individual bias from the students’ perspective. Future work could expand this scope to include bias from the examiners’ perspective, such as investigating inconsistencies among individual officers who perform shortlisting and admission recommendations. This broader assessment could provide a more comprehensive understanding of the biases present in the admission process.
Additionally, our investigation of group bias focused solely on gender. Future research could extend this analysis to include other demographic traits, such as nationality, race, and religion. By doing so, we can explore whether further debiasing efforts lead to improved model performance and more equitable decision-making.
Furthermore, while studying fairness in human decisions and ML predictions, we consistently leverage a pre-trained language model (PLM) i.e. BERT to perform text embeddings. Nonetheless, the PLM may introduce algorithmic bias which we yet to examine.
Lastly, this work did not leverage Large Language Models (LLMs) to perform various tasks. Future studies could assess the effectiveness of using LLMs for debiasing data, potentially enhancing the efficiency and accuracy of the debiasing process.
VII Conclusion
Fairness in decision-making is critical, whether conducted by machine learning models or human experts. Our work first investigates individual fairness, using a consistency score to evaluate decision-making among a diverse group of subjects. Experimental results reveals that ML models such as XGBoost and BiLSTM achieve significantly higher consistency scores than human experts. This finding suggests that ML models can provide more reliable and fair treatment of similar applications, addressing the subjectivity and cognitive biases inherent in human decision-making.
In examining group fairness, we propose a gender-debiasing pipeline that effectively removes gender-specific language from the features without compromising prediction performance. Our experimental results demonstrate that all three models such as XGBoost, BiLSTM and KNN maintain or improve on classification accuracy using debiased data. This validates our hypothesis that fairness and performance can coexist, challenging the assumed trade-off between the two.
Our findings have practical implications for the design of admission systems, suggesting that ML models could serve as valuable decision support tools, particularly in initial screening stages where consistency is crucial. The models could help identify potential inconsistencies in human decisions, prompting additional review in cases where machine and human assessments significantly diverge. Furthermore, the quantification of consistency scores provides a new dimension for evaluating and improving admission processes, beyond traditional accuracy metrics.
Moving forward, we could extend this work by examining biases from the examiners’ perspective and other demographic traits such as nationality, race, and religion. Furthermore, we could investigate potential bias introduced by PLM used for text embedding. Finally, we could explore leveraging Large Language Models (LLMs) for debiasing data to enhance the efficiency and accuracy of the debiasing process.